Example :

In tossing a coin once , Find the probability of events ?

1) The event A is the first tossing is heads

2) The event B is the first tossing is tails

3) The event c is the getting is one at least heads

4) The event c is the getting is one at most tails

--------------------------------

Example :

In tossing a coin once , Find the probability of events ?

1) The event A is the first tossing is heads

2) The event B is the first tossing is tails

3) The event c is the getting is one at least heads

4) The event c is the getting is one at most tails

-----------------

Example :

When you with draw a card from a set of cards number 10 numbered from 1 to 10 .

1)What is the probability of obtaining an initial number or number that can be divided by 3 ?
Answer #1

1) The event A is the first tossing is heads = 1/2

2) The event B is the first tossing is tails = 1/2

3) The event c is the getting is one at least heads=1/2

4) The event c is the getting is one at most tails= P(X=0) + P(X=1) = 1/2+1/2 = 1

-----------------

the probability of obtaining an initial number or number that can be divided by 3 = 3/10 = 0.30